# Resonant excitation of quasinormal modes of black holes

### v1

@@ Abstract
Let us consider the case where continuous variation of $V$ is driven by a single parameter. At each point except the avoided crossing the perturbation theory is valid, and hence $\omega_{i}^{2}$ follow $\mathcal{E}_{i}$. Suppose that the perturbation theory predicts the two modes approach on the same line, for which, denoting $\mathcal{E}_{d}=pe^{i\alpha}$ and $\mathcal{E}_{c}=\omega_{\rm EP}^{2}$, $p$ is a real parameter and $\alpha$ and $\omega_{\rm EP}^{2}$ are constant. We also neglect the variation of $\Delta=qe^{i\beta}$. The single-parameter degree of freedom of $V$ is thus represented by $p$. Denoting $\omega_{\pm}^{2}-\omega_{\rm EP}^{2}=x+iy$, it is straightforward to derive $X^{2}-Y^{2}=q^{2}\sin 2(\alpha-\beta)$, where $X+iY$ is defined by rotation of $x+iy$ by $\pi/4-\alpha$. Unless $\alpha=\beta$, which requires a tuning of an additional parameter, $\omega_{\pm}^{2}-\omega_{\rm EP}^{2}$ follows a hyperbolic avoided crossing, and so does $\omega_{\pm}-\omega_{\rm EP}$ itself, since this occurs at a small region near EP. A simple calculation shows $|\omega_{\pm}^{2}-\omega_{\rm EP}^{2}|\approx({\rm const}+p^{2})^{1/4}$ at the peak when $p\ll 1$, which means $|\omega_{+}-\omega_{-}|^{-1}$ obeys the quarter-power Lorentzian.
@@
On the other hand, QNM eigenstates $\psi_{\pm}$ associated to $\omega_{\pm}$ are given by $\psi_{+}=\psi_{1}+T\psi_{2}$ and $\psi_{-}=-T\psi_{1}+\psi_{2}$, where $T=1/(\delta+\sqrt{\delta^{2}-1})$ is governed by a dimensionless complex parameter $\delta=\sqrt{\mathcal{E}_{d}^{2}/\Delta^{2}+1}$. If the two modes are close and then $\delta\approx 0$ and $\mathcal{A}_{1}^{2}\approx\mathcal{A}_{2}^{2}$ are satisfied, the eigenstates are almost degenerate $\psi_{-}\approx i\psi_{+}$, and the norm squared nearly vanish as $\mathcal{A}^{2}_{\pm}\propto\pm\delta$. These behaviors are nothing but what is expected at EP Özdemir et al. (2019); Wiersig (2020); Parto et al. (2021); Ding et al. (2022). With (6) and (8), we obtain $B_{\pm}\propto\mathcal{A}^{-2}_{\pm}\propto(\omega_{\pm}^{2}-\omega_{\rm EP}^{2})^{-1}$. Hence, the excitation factors are amplified following the inverse of the hyperbola, i.e., the lemniscate of Bernoulli, and the resonance peak obeys the quarter-power Lorentzian.
@@
Conclusions and Discussion. When two QNM frequencies approach on the complex plane, they are anomalously excited. We clarified this resonance phenomenon by developing high-precision numerical analysis and theoretical framework of QNMs in the language of non-Hermitian physics. The resonance of Kerr QNMs not only allows us to decipher a long-standing mystery in black hole physics but also stands as a novel beacon for characterizing black hole spacetime geometry. Deviations from Kerr, whether due to astrophysical effects or correction to general relativity, will manifest in the resonances in QNMs. Our findings thus pave the way for rigorous examinations of black holes and the exploration of new physics in gravity.
Acknowledgments. This work was supported by Japan Society for the Promotion of Science (JSPS) Grants-in-Aid for Scientific Research (KAKENHI) Grant No. JP22K03639.
@@
In the main text we focus on the resonance at avoided crossing in the vicinity of EP of Kerr QNMs. To argue that this phenomenon is a universal feature of QNMs, it is helpful to study a simple toy model away from Kerr QNMs. Indeed, the simplest rectangular barrier potential played an important role in a pioneering work Chandrasekhar and Detweiler (1975) in the early days of QNM research. Along the same line, here we consider a simple toy model
$\left(\frac{d^{2}}{dx^{2}}+\omega^{2}-V(x)\right)\psi=0,$ (9)
@@
$V(x)=\begin{cases}V_{0},&0\leq x\leq d,\\ V_{1},&b\leq x\leq b+d,\\ 0,&{\rm otherwise},\end{cases}$ (10)
with $0\leq d\leq b$ and $V_{0}\geq 0$ and $V_{1}\geq 0$. We define the QNM frequencies by requiring the purely right-/left-going boundary condition, i.e., $\psi\to e^{\pm i\omega x}$, at $x>b+d$ and $x<0$, respectively. We can then calculate the excitation factor by (2) or (6). This toy model allows us to examine the intrinsic behavior of QNMs and excitation factors in a simple and transparent manner.
@@
The QNMs for a similar setup was investigated in Barausse et al. (2014); Cheung et al. (2022); Kyutoku et al. (2023) as the simplest variation of the “flea on the elephant” effect Jona-Lasinio et al. (1981); Graffi et al. (1984); Simon (1985); Santarsiero and Gori (2019), which has been explored recently in the context of the spectral instability Nollert (1996); Barausse et al. (2014); Daghigh et al. (2020); Jaramillo et al. (2021); Qian et al. (2021); Cheung et al. (2022); Jaramillo et al. (2022); Berti et al. (2022); Kyutoku et al. (2023); Cardoso et al. (2024). It has been known that small deformation of potential at far distance causes a drastic change of QNM frequencies while changes in the time-domain ringdown signal shows up only at late time. However, the behavior of the excitation factors has not been clarified.
Let us classify the QNM frequencies for the double rectangular barriers (10) into two categories: The first sequence $\omega^{(1)}_{n}$ $(n=0,1,2,\cdots)$ denotes modes that are smoothly connected to the unperturbed modes with $V_{1}=0$, whereas the second sequence $\omega^{(2)}_{k}$ $(k=0,1,2,\cdots)$ denotes modes that show up when $V_{1}>0$ and correspond to the resonance frequencies for the effective cavity between two barriers. For small $b$, the first sequence modes follow the logarithmic spiral around the unperturbed modes which is consistent with Leung et al. (1997). By using the perturbation theory formula (7), we obtain
$\omega^{(1)}_{n}\simeq\bar{\omega}_{n}-\frac{i\bar{B}_{n}V_{1}\sin\bar{\omega}_{n}d}{\bar{\omega}_{n}}e^{i\bar{\omega}_{n}(2b+d)},$ (11)
@@
$\omega^{(2)}_{k}\simeq\frac{1}{2b}\left[(2k+1)\pi-i\ln\frac{V_{0}}{V_{1}}\right].$ (12)
@@
Thus, as we increase $b$, the second sequence modes accumulate to the origin of the complex frequency plane, during which the second sequence modes (or more generally, migrating modes including modes that were originally belonging to the first sequence) approach to the first sequence modes, which then begin to migrate together with the second sequence modes.
@@
We find that when two QNM frequencies approach each other during the migration, they exhibit the same behavior observed for the Kerr QNMs; The two QNM frequencies repel each other, during which the excitation factors are point-symmetrically amplified. In the left and middle columns of Fig. 4, we depict the trajectories of QNM frequencies and excitation factors with respect to the variation of $b$. In addition to $b$, we control the second parameter $V_{1}$ to see how the avoided crossing and resonance change. The cases with $V_{1}d^{2}=4.30\times 10^{-3}$ and $4.31\times 10^{-3}$ demonstrate sharp resonances, where the trajectories follow the hyperbola and lemniscate. The two cases switch about 90 degrees on the complex plane from each other, which is predicted at the sign flip of the RHS of $X^{2}-Y^{2}=q^{2}\sin 2(\alpha-\beta)$ in the main text. The cases with $V_{1}d^{2}=3.50\times 10^{-3}$ and $5.50\times 10^{-3}$ show mild resonances. The trajectories of the QNM frequencies and excitation factors can be regarded as distorted versions of the hyperbola and lemniscate, respectively. These sharp and mild resonances reproduce what we observed in the Kerr QNM frequencies and excitation factors in the main text.
@@
We also examine the characteristic resonant peak appearing in the difference of the excitation factors and the inverse of the difference of the QNM frequencies. As shown in the right column of Fig. 4, the peaks are well fitted by the quarter-power Lorentzian, same as the Kerr case. For the sharp resonances, the relative difference between any pair of $\Delta B$, $1/\Delta\omega$ (normalized by multiplying a numerical constant), and the quarter-power Lorentzian remains $\mathcal{O}(1)$in the plotted region. The mild resonances are still better fitted by the quarter-power Lorentzian over the Lorentzian itself.
@@
Note also that the resonance phenomenon occurs between the fundamental mode and third overtone of the first sequence. It implies that, while the resonances for Kerr QNMs studied in the main text occur for higher overtones, the resonances involving the fundamental mode and/or lower overtones are in general not prohibited. Indeed, we analyze the double rectangular barriers with various parameter sets and confirm the resonances for various pairs of two modes approaching each other. It suggests that the resonance is a universal feature of QNMs.
@@
Figure 4: Left and middle columns: QNM frequencies (left column) and excitation factors (middle column) for the double rectangular barriers (10). The fundamental mode $\omega^{(1)}_{0}$ (red) and third overtone $\omega^{(1)}_{3}$ (green) of the first sequence are highlighted, while overtones up to the seventh in the first (colored) and second (black) sequences are depicted. We set parameters $V_{0}d^{2}=16$ and $V_{1}d^{2}=(3.50,4.30,4.31,5.50)\times 10^{-3}$ (from first to fourth row) and depict trajectories with respect to the variation of $b/d$ in the range $1\leq b/d\leq 15$ with arrows indicating the direction of increasing $b/d$. The logarithmic spiral (black dotted), hyperbola (black dashed), and lemniscate of Bernoulli (black dashed) approximate the trajectories well. Right column: $|B^{(1)}_{0}-B^{(1)}_{3}|$ (blue solid) and $1/|\omega^{(1)}_{0}-\omega^{(1)}_{3}|$ (orange dashed), the latter of which is multiplied by a numerical constant to match the maxima. The peak can be well fitted by a quarter power of Lorentzian (black dotted) rather than the Lorentzian itself (gray dotted).

### v2

@@ Abstract
Let us consider the case where continuous variation of $V$ is driven by a single parameter. At each point except the avoided crossing the perturbation theory is valid, and hence $\omega_{n_{i}}^{2}$ follow $\mathcal{E}_{n_{i}}$. Suppose that the perturbation theory predicts that the two modes approach each other along the same line, for which, denoting $\mathcal{E}_{d}=pe^{i\alpha}$ and $\mathcal{E}_{c}=\omega_{\rm EP}^{2}$, $p$ is a real parameter and $\alpha$ and $\omega_{\rm EP}^{2}$ are constant. We also neglect the variation of $\Delta=qe^{i\beta}$. The single-parameter degree of freedom of $V$ is thus represented by $p$. Denoting $\omega_{\pm}^{2}-\omega_{\rm EP}^{2}=x+iy$, it is straightforward to derive $X^{2}-Y^{2}=q^{2}\sin 2(\alpha-\beta)$, where $X+iY$ is defined by a rotation of $x+iy$ by $\pi/4-\alpha$. Unless $\alpha=\beta$, which would require tuning an additional parameter, $\omega_{\pm}^{2}-\omega_{\rm EP}^{2}$ follows a hyperbolic avoided crossing, and so does $\omega_{\pm}-\omega_{\rm EP}$ itself, since this occurs at a small region near the EP. A simple calculation shows $|\omega_{\pm}^{2}-\omega_{\rm EP}^{2}|\approx({\rm const}+p^{2})^{1/4}$ at the peak when $p\ll 1$, meaning that $|\omega_{+}-\omega_{-}|^{-1}$ obeys the quarter-power Lorentzian.
@@
On the other hand, QNM eigenstates $\psi_{\pm}$ associated with $\omega_{\pm}$ are given by $\psi_{+}=\psi_{n_{1}}+T\psi_{n_{2}}$ and $\psi_{-}=-T\psi_{n_{1}}+\psi_{n_{2}}$, where $T=1/(\delta+\sqrt{\delta^{2}-1})$ is governed by a dimensionless complex parameter $\delta=\sqrt{\mathcal{E}_{d}^{2}/\Delta^{2}+1}$. If the two modes are close and then $\delta\approx 0$ and $\mathcal{A}_{n_{1}}^{2}\approx\mathcal{A}_{n_{2}}^{2}$ are satisfied, the eigenstates are almost degenerate $\psi_{-}\approx i\psi_{+}$, and the norm squared nearly vanishes as $\mathcal{A}^{2}_{\pm}\propto\pm\delta$. These behaviors are nothing but what is expected at the EP [47, 48, 49, 50]. With Eqs. (6) and (8), we obtain $B_{\pm}\propto\mathcal{A}^{-2}_{\pm}\propto(\omega_{\pm}^{2}-\omega_{\rm EP}^{2})^{-1}$. Hence, the excitation factors are amplified following the inverse of the hyperbola, i.e., the lemniscate of Bernoulli, and the resonance peak obeys the quarter-power Lorentzian.
@@
Conclusions and Discussion—When two QNM frequencies approach each other on the complex plane, they exhibit anomalously amplified excitations. We have elucidated this resonance phenomenon near EPs by developing high-precision numerical analysis and theoretical framework of QNMs in the language of non-Hermitian physics. The resonance of QNMs not only allows us to decipher the long-standing mystery in black hole physics but also stands as a novel beacon in black hole spectroscopy. Deviations from Kerr black holes, whether due to astrophysical effects or corrections to general relativity, will manifest through the resonances and be imprinted in the ringdown GWs.
Notably, similar enhancement effects around EPs in optics have already been employed to realize optical sensors with enhanced performance [48, 49]. Additionally, it is well known that the Mikheyev-Smirnov-Wolfenstein resonance at the avoided crossing of neutrino eigenstates in matter plays a key role in solar neutrino spectroscopy [43, 44, 45]. With high signal-to-noise ratio ringdown GWs from black holes with various spins expected in the near future, the resonant excitation of QNMs opens the door to rigorous examinations of black holes and the discovery of new physics in gravity.
Acknowledgments—This work was supported by JSPS KAKENHI Grant No. JP22K03639.
Data availability—The data that support the findings of this article are openly available [62].
@@
In the main text we focus on the resonance at avoided crossing in the vicinity of EP of Kerr QNMs. To demonstrate that this phenomenon is a universal feature of QNMs, it is helpful to study a simple toy model apart from Kerr QNMs. Indeed, the simplest rectangular barrier potential played an important role in a pioneering work [11] in the early days of QNM research. Along the same line, here we consider a simple toy model
$\left(\frac{d^{2}}{dx^{2}}+\omega^{2}-V(x)\right)\psi=0,$ (S1)
@@
$V(x)=\begin{cases}V_{0},&0\leq x\leq d,\\ V_{1},&b\leq x\leq b+d,\\ 0,&{\rm otherwise},\end{cases}$ (S2)
with $0\leq d\leq b$ and $V_{0}\geq 0$ and $V_{1}\geq 0$. We define the QNM frequencies by requiring the purely right-/left-going boundary condition, i.e., $\psi\to e^{\pm i\omega x}$, at $x>b+d$ and $x<0$, respectively. We can then calculate the excitation factor by (2) or (6). This toy model allows us to examine the intrinsic behavior of QNM frequencies and excitation factors in a simple and transparent manner.
@@
The QNMs for a similar setup was investigated in [83, 84, 85] as the simplest variation of the “flea on the elephant” effect [86, 87, 88, 89], which has been explored recently in the context of the spectral instability [90, 83, 91, 104, 92, 84, 93, 94, 85, 95]. It has been known that small deformation of potential at far distance causes a drastic change of QNM frequencies while changes in the time-domain ringdown signal show up only at late time. However, the behavior of the excitation factors has not been clarified.
Let us classify the QNM frequencies for the double rectangular barriers (S2) into two categories: The first sequence $\omega^{(1)}_{n}$ $(n=0,1,2,\cdots)$ denotes modes that are smoothly connected to the unperturbed modes with $V_{1}=0$. The second sequence $\omega^{(2)}_{k}$ $(k=0,1,2,\cdots)$ denotes modes that show up when $V_{1}>0$ and correspond to the resonance frequencies for the effective cavity between the two barriers. The labeling of overtone indices $n$ and $k$ is determined based on their decay times at $V_{1}\to 0$.
For small $b$, the first sequence modes follow the logarithmic spiral around the unperturbed modes which is consistent with [100]. By using the perturbation theory formula (7), we obtain
$\omega^{(1)}_{n}\simeq\bar{\omega}_{n}-\frac{i\bar{B}_{n}V_{1}\sin\bar{\omega}_{n}d}{\bar{\omega}_{n}}e^{i\bar{\omega}_{n}(2b+d)},$ (S3)
@@
$\omega^{(2)}_{k}\simeq\frac{1}{2b}\left[(2k+1)\pi-i\ln\frac{V_{0}}{V_{1}}\right].$ (S4)
@@
Thus, as we increase $b$, the second sequence modes accumulate to the origin of the complex frequency plane, during which the second sequence modes (or migrating modes more generally, including modes that were originally belonging to the first sequence) approach to the first sequence modes, which then begin to migrate together with the migrating modes.
@@
We find that when two QNM frequencies approach each other during the migration, they exhibit the same behavior observed for the Kerr QNMs: The two QNM frequencies repel each other, during which the excitation factors are point-symmetrically amplified. In the left and middle columns of Fig. S1, we depict the trajectories of QNM frequencies and excitation factors with respect to the variation of $b$. In addition to $b$, we control the second parameter $V_{1}$ to see how the avoided crossing and resonance change. The cases with $V_{1}d^{2}=4.30\times 10^{-3}$ and $4.31\times 10^{-3}$ demonstrate sharp resonances, where the trajectories follow the hyperbola and lemniscate. The two cases exhibit a 90-degree switch in the complex plane, which is predicted at the sign flip of the RHS of $X^{2}-Y^{2}=q^{2}\sin 2(\alpha-\beta)$ in the main text. The cases with $V_{1}d^{2}=3.50\times 10^{-3}$ and $5.50\times 10^{-3}$ demonstrate mild resonances, where the trajectories of the QNM frequencies and excitation factors can be regarded as distorted versions of the hyperbola and lemniscate, respectively. These sharp and mild resonances reproduce what we observed in the Kerr QNM frequencies and excitation factors in the main text.
@@
We also examine the characteristic resonant peak appearing in the difference of the excitation factors and the inverse of the difference of the QNM frequencies. As shown in the right column of Fig. S1, the peaks are well fitted by the quarter-power Lorentzian, same as the Kerr case. For the sharp resonances, the relative difference between any pair of $\Delta B$, $1/\Delta\omega$ (normalized by multiplying a numerical constant), and the quarter-power Lorentzian remains $\mathcal{O}(1)$in the plotted region. The mild resonances are still better fitted by the quarter-power Lorentzian over the Lorentzian itself.
@@
Note also that the resonance phenomenon occurs between the fundamental mode and third overtone of the first sequence. This reveals that, while the resonances for Kerr QNMs studied in the main text primarily involve higher overtones, the resonances involving the fundamental mode and/or lower overtones are in general not prohibited. Indeed, we analyze the double rectangular barriers across various parameter sets and confirm the resonances for various pairs of modes approaching each other. This supports that the resonance is a universal feature of QNMs.
@@
Figure S1: Left and middle columns: QNM frequencies (left column) and excitation factors (middle column) for the double rectangular barriers (S2). The fundamental mode $\omega^{(1)}_{0}$ (red) and third overtone $\omega^{(1)}_{3}$ (green) of the first sequence are highlighted, while overtones up to the seventh in the first (colored) and second (black) sequences are depicted. We set parameters $V_{0}d^{2}=16$ and $V_{1}d^{2}=(3.50,4.30,4.31,5.50)\times 10^{-3}$ (from first to fourth row) and depict trajectories with respect to the variation of $b/d$ in the range $1\leq b/d\leq 15$ with arrows indicating the direction of increasing $b/d$. The logarithmic spiral (black dotted), hyperbola (black dashed), and lemniscate of Bernoulli (black dashed) approximate the trajectories well. Right column: $|B^{(1)}_{0}-B^{(1)}_{3}|$ (blue solid) and $1/|\omega^{(1)}_{0}-\omega^{(1)}_{3}|$ (orange dashed), the latter of which is multiplied by a numerical constant to match the maxima. The peak can be well fitted by a quarter power of Lorentzian (black dotted) rather than the Lorentzian itself (gray dotted).